OpenStudy (anonymous):

4x-5/10 ≤-4 or 4x-5/10≥7 solve the compound inequality

OpenStudy (anonymous):

ok it is OR yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{4x-5}{10} \leq -4\]

OpenStudy (saifoo.khan):

is the whole thing over thing?

OpenStudy (anonymous):

it has to have two answers .

OpenStudy (anonymous):

\[4x-5\leq -40\] \[4x\leq -35\] \[x\leq-\frac{35}{4}\]

OpenStudy (anonymous):

looks same as last one.

OpenStudy (anonymous):

that is one inequality. the other is \[\frac{4x-5}{10}\geq7\] \[4x-5\geq 70\] \[4x\geq 75\] \[x\geq \frac{75}{4}\]

OpenStudy (anonymous):

two, i repeat two intervals \[(-\infty, -\frac{35}{4}]\cup [\frac{75}{4},\infty)\]

OpenStudy (anonymous):

no it should look like this this is the way the answer reads {x│x < ____ or x> ____}

OpenStudy (anonymous):

Mathematica agrees with satellite73\[\text{Reduce}[(4x-5)/10\leq -4 \|( 4x-5)/10\geq 7]\to x\leq -\frac{35}{4}\|x\geq \frac{75}{4} \]

OpenStudy (anonymous):

no it should look like this this is the way the answer reads {x│x < ____ or x> ____}

OpenStudy (anonymous):

ok you are writing as in inequality

OpenStudy (anonymous):

that answer shows incorrect.

OpenStudy (anonymous):

\[{x|x \le-35/4 or x \ge75/4}\]

OpenStudy (anonymous):

that is \( \{ x \; | \; x \le \frac{-35}4 \text{ or } x \ge \frac{75}4 \} \)